# FC-Attack: Jailbreaking Large Vision-Language Models via Auto-Generated Flowcharts

### v1

@@ Introduction
We then explore the impact of different factors in $\mathsf{FC\text{-}Attack}$ on jailbreak performance, including the different types of user queries, the number of descriptions in flowcharts, and the font styles used in flowcharts.
Different Types of User Query. We investigate whether the content in flowcharts, when directly input as text, can lead to the jailbroken of LVLMs. The flowchart content consists of two parts: harmful query from AdvBench and the step descriptions generated by the generator based on this query.
As shown in Figure 4 when using only the harmful query (text) as input, we observe very low ASR. The ASR is $0\%$ on GPT-4o mini, GPT-4o, Claude-3.5, Qwen2-VL, and InternVL-2.5, and only $2\%$ and $6\%$ on Gemini-1.5 and Llava-Next, respectively. This indicates that the textual modality of these LVLMs has relatively robust defenses against such inputs. However, when both the harmful query and the step descriptions are input as text, the ASR increases to $36\%$ on Gemini-1.5, and to $16\%$ and $6\%$ on Llava-Next and InternVL-2.5, respectively, while remaining at $0\%$ on the other models. When this information is converted into a flowchart and only a benign textual prompt is provided, the ASR on these models improves significantly. This demonstrates that the defenses of LVLMs in the visual modality have noticeable weaknesses compared with the language modality.
Table 4: Comparison of ASR for different defense methods across various LVLMs.
@@
However, more descriptions are not always better. For example, for the Gemini-1.5 model, the Vertical flowcharts achieve their highest ASR of $82\%$ at “3 steps” but slightly drop to $80\%$ at 5 steps and full descriptions. A similar trend is observed in Horizontal and S-shaped flowcharts, where ASR reaches $88\%$ and $86\%$ at “5 steps” but decreases to $76\%$ and $74\%$, respectively, at full descriptions. This phenomenon may be related to the resolution processing capability of LVLMs. When the number of descriptions increases to full, the descriptions may include redundant information, which could negatively impact the model’s performance.
Font Styles in Flowcharts. To investigate whether different font styles in flowcharts affect the effectiveness of jailbreak attacks, we select five fonts from Google Fonts that are relatively difficult for humans to read: Creepster, Fruktur Italic, Pacifico, Shojumaru, and UnifrakturMaguntia (the font style examples are shown in Figure 5). Table 3 shows the results of $\mathsf{FC\text{-}Attack}$ (Ensemble). We observe that different font styles can significantly impact the ASR. For example, on GPT-4o mini, the ASR increases across all font styles compared to the original, with Shojumaru font achieving the highest ASR of $20\%$. Similarly, on Claude-3.5, the Pacifico font achieves the highest ASR of $28\%$, which is a substantial improvement compared to the original ASR of $4\%$. For Gemini-1.5, the ASR reaches $98\%$ with the Fruktur font, while Llava-Next achieves $94\%$ with the Shojumaru font. InternVL-2.5 also shows a $6\%$ increase in ASR with the Pacifico font, reaching $96\%$. These findings further highlight the need to consider the impact of different font styles when designing defenses.
Table 5: Performance comparison on the Benign Dataset (MM-vet [35]) under Adashield-S (Ada-S) and Adashield-A (Ada-A) is presented across various LVLMs, reporting scores on six core vision-language capabilities: Recognize (Rec), OCR, Knowledge (Know), Generation (Gen), Spatial (Spat), and Math.
@@
As shown in Table 4, we test four methods to defend against $\mathsf{FC\text{-}Attack}$, where “Original” represents the results of $\mathsf{FC\text{-}Attack}$ (Ensemble) with an average ASR of $58.6\%$. Using Llama-Guard3-V and JailGuard to detect whether the input is harmful reduced the ASR to $51.7\%$ and $51.4\%$, respectively. This might be because the information in flowcharts is primarily presented in textual form, while these detection methods are better at identifying information in visual content, making them less effective at detecting textual information. AdaShield-S and AdaShield-A reduce the average ASR to $6.9\%$ and $1.7\%$, respectively, by adding defense prompts in the input, showing more effective defense performance.
However, these two methods also lead to a decline in LVLMs performance on benign datasets. We conduct tests on MM-Vet [35] to evaluate the important factor of “over-defensiveness” on benign datasets, which is an evaluation benchmark that contains complex multimodal tasks for LVLMs. As shown in Table 5, the model’s utility decreases on benign data when using AdaShield-S and AdaShield-A, indicating a future direction for defense development.
@@ Conclusion
In this paper, we propose $\mathsf{FC\text{-}Attack}$, which leverages auto-generated flowcharts to jailbreak LVLMs. Experimental results demonstrate that $\mathsf{FC\text{-}Attack}$ achieves higher ASR in both open-source and production LVLMs compared to other jailbreak attack methods. Additionally, we investigate the factors influencing $\mathsf{FC\text{-}Attack}$, including different types of user queries, the number of steps in flowcharts, and font styles in flowcharts, gaining insights into the aspects that affect ASR. Finally, we explore several defense strategies and demonstrate that the AdaShield-A method can effectively mitigate $\mathsf{FC\text{-}Attack}$ but with the cost of utility drop.
@@
This paper presents a method, $\mathsf{FC\text{-}Attack}$, for jailbreaking LVLMs using harmful flowcharts. As long as the adversary obtains a harmful flowchart, they can jailbreak LVLMs with minimal resources. Therefore, it is essential to systematically identify the factors that influence the attack success rate and offer potential defense strategies to model providers. Throughout this research, we adhere to ethical guidelines by refraining from publicly distributing harmful flowcharts and harmful responses on the internet before informing service providers of the risks. Prior to submitting the paper, we have already sent a warning e-mail to the model providers about the dangers of flowchart-based jailbreak attacks on LVLMs and provided them with the flowcharts generated in our experiments for vulnerability mitigation.
Appendix A Introduction of LVLMs in this paper
In this section, we introduce the LVLMs used in this paper.
  • Llava-Next (January 2024) is an open-source LVLM released by the University of Wisconsin-Madison, which builds upon the Llava-1.5 model [21] with multiple improvements [22]. It enhances capabilities in visual reasoning, optical character recognition, and world knowledge. Besides, Llava-Next increases the input image resolution to a maximum of $672\times 672$ pixels and supports various aspect ratios to capture more visual details ($336\times 1344$ and $1344\times 336$).
@@ Appendix A Introduction of LVLMs in this paper
  • Gemini-1.5 (February 2024) is a production-grade LVLM developed by Google, based on the Mixture-of-Experts architecture [12]. For Gemini-1.5, larger images will be scaled down to the maximum resolution of $3072\times 3072$, and smaller images will be scaled up to $768\times 768$ pixels. Reducing the image size will not improve the performance of higher-resolution images.
@@
  • GPT-4o and GPT-4o Mini are popular production-grade LVLMs developed by OpenAI [14, 25]. GPT-4o Mini is a compact version of GPT-4o, designed for improved cost-efficiency. Both models excel in handling complex visual and language understanding tasks.
  • InternVL-2.5 (June 2024) [10] is an open-source LVLM that ranks first in full-scale open-source multimodal performance. In terms of multimodal long-chain reasoning, it achieves a breakthrough of $70\%$ in the expert-level multidisciplinary knowledge reasoning benchmark MMMU [36], and the general capabilities are significantly enhanced.
@@ Appendix C Generator Setting
(a) Flowchart Example-Horizontal: Left-to-right flow
@@
(c) Flowchart Example-Vertical: Top-to-bottom flow
Figure A1: Flowchart Examples: Various flow directions

### v2

@@ Introduction
We then explore the impact of different factors in $\mathsf{FC\text{-}Attack}$ on jailbreak performance, including the different types of user queries, the number of descriptions, and the font styles used in flowcharts.
Different Types of User Query. We investigate whether the content in flowcharts, when directly input as text, can lead to the jailbroken of MLLMs. The flowchart content consists of two parts: harmful query from AdvBench and the step descriptions generated by the generator based on this query.
As shown in Figure 3 when using only the harmful query (text) as input, we observe very low ASR. The ASR is $0\%$ on GPT-4o mini, GPT-4o, Claude-3.5, Qwen2-VL, and InternVL-2.5, and only $2\%$ and $6\%$ on Gemini-1.5 and Llava-Next, respectively. This indicates that the textual modality of these MLLMs has relatively robust defenses against such inputs. However, when both the harmful query and the step descriptions are input as text, the ASR increases to $36\%$ on Gemini-1.5, and to $16\%$ and $6\%$ on Llava-Next and InternVL-2.5, respectively, while remaining at $0\%$ on the other models. When this information is converted into a flowchart and only a benign textual prompt is provided, the ASR on these models improves significantly. This demonstrates that the defenses of MLLMs in the visual modality have noticeable weaknesses compared with the language modality.
Table 5: Comparison of ASR for different defense methods across various MLLMs.
@@
However, more descriptions are not always better. For example, for the Gemini-1.5 model, the Vertical flowcharts achieve their highest ASR of $82\%$ at “3 steps” but slightly drop to $80\%$ at 5 steps and full descriptions. A similar trend is observed in Horizontal and S-shaped flowcharts, where ASR reaches $88\%$ and $86\%$ at “5 steps” but decreases to $76\%$ and $74\%$, respectively, at full descriptions. This phenomenon may be related to the resolution processing capability of MLLMs. When the number of descriptions increases to full, the descriptions may include redundant information, which could negatively impact the model’s performance.
Font Styles in Flowcharts. To investigate whether different font styles in flowcharts affect the effectiveness of jailbreak attacks, we select five fonts from Google Fonts that are relatively difficult for humans to read: Creepster, Fruktur Italic, Pacifico, Shojumaru, and UnifrakturMaguntia (the font style examples are shown in Figure A3). Table 3 shows the results of $\mathsf{FC\text{-}Attack}$ (Ensemble). We observe that different font styles can significantly impact the ASR. For example, on GPT-4o mini, the ASR increases across all font styles compared to the original, with Shojumaru font achieving the highest ASR of $20\%$. Similarly, on Claude-3.5, the Pacifico font achieves the highest ASR of $28\%$, which is a substantial improvement compared to the original ASR of $4\%$. For Gemini-1.5, the ASR reaches $98\%$ with the Fruktur font, while Llava-Next achieves $94\%$ with the Shojumaru font. InternVL-2.5 also shows a $6\%$ increase in ASR with the Pacifico font, reaching $96\%$. These findings further highlight the need to consider the impact of different font styles when designing defenses.
Table 6: MLLM performance on the Benign MM-Vet dataset [43] under Adashield-S (Ada-S) and Adashield-A (Ada-A), covering six core tasks: Recognize (Rec), OCR, Knowledge (Know), Generation (Gen), Spatial (Spat), and Math.
@@
Effect of Flowchart Structure. To explore the impact of graphical structure elements on the jailbreak effect. We conduct experiments with Qwen2-VL using four different flowchart designs: (1) an enhanced FigStep flowchart where each step incorporates step descriptions generated by $\mathsf{FC\text{-}Attack}$; (2) Plain Text structure that only retains text without any graphical elements in the flowchart; (3) Text with Box structure that encapsulates each step in boxes but omits directional arrows; and (4) our complete $\mathsf{FC\text{-}Attack}$ implementation featuring both boxes surrounding step descriptions and arrows indicating the progression between steps. Table A1 shows the results of four flowchart image structures. We notice that the ASR of the FigStep method is 34%, that of Plain Text is 32%, that of Text with Box is 50%, and that of $\mathsf{FC\text{-}Attack}$ is 88%. It is noted that the addition of box elements improves ASR by 18%, while the introduction of directional arrows connecting these boxes further improves it by 38%. These findings reveal the contribution of the graphical structural elements of the flowchart to improving the jailbreak effect.
@@
We consider four defenses (shown in Table 5), where “Original” represents the results of $\mathsf{FC\text{-}Attack}$ (Ensemble) with an average ASR of $58.6\%$. Using Llama-Guard3-V and JailGuard to detect whether the input is harmful reduced the ASR to $51.7\%$ and $51.4\%$, respectively. The limited effectiveness may stem from flowcharts being primarily text-based, whereas the detection methods are more suited to visual content. AdaShield-S and AdaShield-A reduce the average ASR to $6.9\%$ and $1.7\%$, showing more effective defense performance. However, these two methods also lead to a decline in MLLMs performance on benign datasets. We conduct tests on MM-Vet [43] to evaluate the important factor of “over-defensiveness” on benign datasets, which is an evaluation benchmark that contains complex multimodal tasks for MLLMs. As shown in Table 6, the model’s utility decreases on benign data when using AdaShield-S and AdaShield-A, indicating a future direction for defense development.
@@ Conclusion
In this paper, we propose $\mathsf{FC\text{-}Attack}$, which leverages auto-generated flowcharts to jailbreak MLLMs. Experimental results demonstrate that $\mathsf{FC\text{-}Attack}$ achieves higher ASR in both open-source and production MLLMs compared to other jailbreak attacks. Additionally, we investigate the factors influencing $\mathsf{FC\text{-}Attack}$, including different types of user queries, the number of steps in flowcharts, and font styles in flowcharts, gaining insights into the aspects that affect ASR. Finally, we explore several defense strategies and demonstrate that the AdaShield-A method can effectively mitigate $\mathsf{FC\text{-}Attack}$, but with the cost of utility drop.
Limitations
Our work proposes a novel jailbreak attack on MLLMs via images and videos. However, several limitations remain:
  • Limited language scope: In this study, we only consider jailbreak attacks conducted in English, as it is the most widely used global language. In future work, we plan to explore jailbreak performance in other languages, such as Japanese, Spanish, and Chinese.
  • Limited model coverage: This work evaluates only 10 representative MLLMs. Future studies can expand this analysis to include more and newer models as they emerge.
  • Lack of variation in generation parameters: We used a fixed set of generation parameters (e.g., temperature) throughout our experiments. We did not investigate how different decoding settings might affect the success of jailbreak attacks. We plan to include such analyses in future work.
@@
This paper presents a method, $\mathsf{FC\text{-}Attack}$, for jailbreaking MLLMs using harmful flowcharts. As long as the adversary obtains a harmful flowchart, they can jailbreak MLLMs with minimal resources. Therefore, it is essential to systematically identify the factors that influence the attack success rate and offer potential defense strategies to model providers. Throughout this research, we adhere to ethical guidelines by refraining from publicly distributing harmful flowcharts and harmful responses on the internet before informing service providers of the risks. Prior to submitting the paper, we have already sent a warning e-mail to the model providers about the dangers of flowchart-based jailbreak attacks on MLLMs and provided them with the flowcharts generated in our experiments for vulnerability mitigation. We will release our dataset under the Apache 2.0 License.
Appendix A Introduction of MLLMs in this paper
In this section, we introduce the MLLMs used in this paper.
  • Llava-Next (January 2024) is an open-source MLLM released by the University of Wisconsin-Madison, which builds upon the Llava-1.5 model [25] with multiple improvements [26]. It enhances capabilities in visual reasoning, optical character recognition, and world knowledge. Besides, Llava-Next increases the input image resolution to a maximum of $672\times 672$ pixels and supports various aspect ratios to capture more visual details ($336\times 1344$ and $1344\times 336$).
@@ Appendix A Introduction of MLLMs in this paper
  • Gemini-1.5 (February 2024) is a production-grade MLLM developed by Google, based on the Mixture-of-Experts architecture [15]. For Gemini-1.5, larger images will be scaled down to the maximum resolution of $3072\times 3072$, and smaller images will be scaled up to $768\times 768$ pixels. Reducing the image size will not improve the performance of higher-resolution images.
@@
  • GPT-4o and GPT-4o Mini are popular production-grade MLLMs developed by OpenAI [18, 29]. GPT-4o Mini is a compact version of GPT-4o, designed for improved cost-efficiency. Both models excel in handling complex visual and language understanding tasks.
  • InternVL-2.5 (June 2024) [13] is an open-source MLLM that ranks first in full-scale open-source multimodal performance. In terms of multimodal long-chain reasoning, it achieves a breakthrough of $70\%$ in the expert-level multidisciplinary knowledge reasoning benchmark MMMU [44], and the general capabilities are significantly enhanced.
  • Qwen-vl-max (January 2024) is the most powerful large-scale visual language model developed by the Alibaba team [5]. Compared with the enhanced version, the model has made further improvements in visual reasoning and the ability to follow instructions, providing a higher level of visual perception and cognitive understanding. It provides the best performance on a wider range of complex tasks, can handle a variety of visual understanding challenges, and demonstrates excellent visual analysis capabilities.
  • Qwen2.5-Omni (March 2025) is the new flagship end-to-end multimodal model in the Qwen series [41]. It is designed for comprehensive multimodal perception and can seamlessly handle multiple input forms, including text, images, audio, and video. In addition to its powerful multimodal understanding capabilities, Qwen2.5-Omni can also provide real-time streaming responses through text generation and natural speech synthesis.
  • LLaVA-Video (October 2024) is an extended version of video processing based on the LLaVA-Next model, which inherits the enhanced capabilities of LLaVA-Next in visual reasoning, optical character recognition, and world knowledge, while expanding the processing capabilities of video content [48]. The model supports high-resolution input (up to 672×672 pixels) and various aspect ratios (336×1344 and 1344×336) to capture more visual details, making it perform well in video understanding and analysis.
@@ Appendix C Generator Setting
Appendix D Flowchart Experiment Performance
Figure A1: Successful jailbreak instance on Gemini-1.5 using $\mathsf{FC\text{-}Attack}$.
Figure A2: ASR under different prompts against MLLMs via videos.
Figure A3: Different styles of fonts in flowcharts (“1 step”).
(a) Flowchart Example-Horizontal: Left-to-right flow.
@@ Appendix D Flowchart Experiment Performance
(c) Flowchart Example-Vertical: Top-to-bottom flow.
Figure A4: Flowchart Examples: Various flow directions.
Table A1: Performances of ASR for different flowchart structures on the Qwen-2-VL.
Method $\mathsf{FC\text{-}Attack}$ Plain Text Text with Box Figstep
ASR (%) 88 32 50 34